OpenStudy (jbbadsunsaregrreat):

Soo confused on how to factor this: (𝑥) = 𝑥^3 + 6𝑥^2 − 13𝑥 − 42

TheSmartOne (thesmartone):

Is that supposed to be f(x) = x^3 + 6x^2 − 13x − 42 or x = x^3 + 6x^2 − 13x − 42

OpenStudy (jbbadsunsaregrreat):

@TheSmartOne The 2nd one.

TheSmartOne (thesmartone):

Subtract x on both sides first

OpenStudy (jbbadsunsaregrreat):

@TheSmartOne So It would be like 0=x^3+6x^2-12x-42

TheSmartOne (thesmartone):

-13x - x = -x(13 + 1) = -x(14) = -14x So, it should be -14x not -12x. Do you understand? :)

OpenStudy (jbbadsunsaregrreat):

@TheSmartOne Ohh Okay! Yeah I understand. Thanks!

TheSmartOne (thesmartone):

Now, we should have: x^3 + 6x^2 - 14x - 42 = 0 And it cannot be factored

OpenStudy (daniel.ohearn1):

Can't is such a brash word

OpenStudy (daniel.ohearn1):

From the beginning x^3+6x^2-13x-42 Try picking out a factor of 42 and incorporating x

OpenStudy (daniel.ohearn1):

Clue: try x+2

OpenStudy (daniel.ohearn1):

Just about anything can be factored.

Directrix (directrix):

f(x) = x^3 + 6x^2 − 13x − 42 does factor over the set of Integers.

Directrix (directrix):

x^3 + 6x^2 − 13x − 42 = (x + 7) * (x^2 - x - 6) = (x + 7) * (x - 3) * (x + 2) @JBBADsunsaregrreat Look at this and see if you think it is correct.

TheSmartOne (thesmartone):

That's why I asked in my first reply if he meant x or f(x). :P

OpenStudy (jbbadsunsaregrreat):

@Directrix Yeah I think that is correct! Thanks so much I was getting really frustrated haha

OpenStudy (jbbadsunsaregrreat):

@Directrix If you can, can you show me how to start the problem to come to that answer?

OpenStudy (jbbadsunsaregrreat):

@TheSmartOne Sorry must've made a mistake!

Directrix (directrix):

@JBBADsunsaregrreat When you copied and pasted the problem, the lead "f" that goes with the (x) to be f(x) was deleted. The posted problem looked like this: x = x^3 + 6x^2 − 13x − 42 That is different from f(x) = x^3 + 6x^2 − 13x − 42. Always look back and check the problem you are about to post after the copy and paste procedure.

Directrix (directrix):

@JBBADsunsaregrreat I began by using the Rational Root Theorem.